OpenStudy (arctic_fox):

Choose the correct simplification of the expression a6bc/a4b2?

OpenStudy (proheiper):

:/ another pphalke

OpenStudy (arctic_fox):

what?

OpenStudy (proheiper):

k so

OpenStudy (proheiper):

want me to show you how to get the answer

OpenStudy (proheiper):

or teach you how to do the problem

OpenStudy (arctic_fox):

im not great with the Division of Monomials so yes

OpenStudy (proheiper):

what do you mean "yes"

OpenStudy (proheiper):

that wasnt a "yes" question

OpenStudy (proheiper):

choose one and ill do it

OpenStudy (arctic_fox):

yes id like you to show me how to get the answer, please

OpenStudy (proheiper):

ok so it say "Choose the correct simplification" does that mean its multiple choice

OpenStudy (arctic_fox):

oh yea

OpenStudy (proheiper):

what are the options so i can show you how to eliminate the wrong ones

OpenStudy (proheiper):

...

OpenStudy (arctic_fox):

answers \[\frac{ a2 }{ b }\] a10b5 \[\frac{ a10 }{ b5 }\] a2b

OpenStudy (arctic_fox):

the numbers are exponents

OpenStudy (proheiper):

wait did you write the problem right? a6bc/a4b2

OpenStudy (arctic_fox):

a6b3/a4b2

OpenStudy (arctic_fox):

sorry

OpenStudy (proheiper):

hm

OpenStudy (arctic_fox):

i messed up again ughhh one sec

OpenStudy (arctic_fox):

\[\frac{ a6b3 }{ a4b2}\]

OpenStudy (proheiper):

:) ok

OpenStudy (arctic_fox):

better???

OpenStudy (arctic_fox):

so are you able to help?

OpenStudy (proheiper):

yes

OpenStudy (proheiper):

im kinda stuck too i keep getting 144a2b2

OpenStudy (arctic_fox):

i hate this common core bull

OpenStudy (proheiper):

same

OpenStudy (mathstudent55):

\(\large \dfrac{a^6b^3}{a^4b^2}\) When you divide powers with the same base, subtract the exponents. \(\large \dfrac{a^6b^3}{a^4b^2} = a^{6 - 4} b^{3 - 2} = a^2b^1 = a^2 b\)